# The Origin of Quantum Mechanical Statistics: Some Insights from the Research on Human Language

### v1

@@ 1 Introduction
The foundations of thermodynamics, as prescribed by classical statistical mechanics, rely on the hypothesis that, at a microscopic level, physical systems, or ‘entities’, behave (i) randomly and (ii) independently from each other (Huang,, 1987). A thorough analysis of the historical development of quantum mechanics (QM) reveals that the hypothesis of statistical independence started to be questioned during early quantum theory (1900–1925) and led to the birth of quantum mechanical statistics, in particular, Bose–Einstein statistics (Klein,, 1961; Howard,, 1990; Darrigol,, 1991; Monaldi,, 2009; Perez & Sauer,, 2010; Gorroochurn,, 2018). However, this was done in a way that, in our opinion, is neither straightforward nor completely understood yet.
@@
Historians and philosophers of science have carefully reconstructed early quantum theory through articles and personal correspondence and generally agree that Einstein’s 1905 article on the photoelectric effect (Einstein,, 1905) can be considered as the first genuine article on the quantum nature of physical entities, more than Planck’s 1900 article on the black body radiation (Planck,, 1900). Indeed, Einstein was much more explicit about ‘quantization’ than Planck’s original intention, and he directly assumed the existence of what he called ‘quanta of light’, which we now call ‘photons’ (Klein,, 1961). At the time, Ehrenfest, a student of Boltzmann and close friend of Einstein, was in favour of his idea of quanta of light as entities in their own right, an idea that was not shared by most of the scientific community.
@@
Both Wien and Planck had followed the path of Boltzmann’s thermodynamics, at a time when the so-called ‘ultraviolet catastrophe’ of the classical theory of radiation was still not known. [^3] The expression ‘ultraviolet catastrophe’ was introduced by Ehrenfest only in 1911 (Ehrenfest,, 1911). However, both Einstein and Ehrenfest identified a fundamental problem in Planck’s law, namely, the latter seemed to violate the request of statistical independence of the micro-states of different photons. This, though, is a feature of the Boltzmann view of statistical mechanics, shared by all protagonists, Einstein, Ehrenfest, but also Planck, which also follows as a consequence of the Maxwell–Boltzmann distribution defining the governing statistics. The latter result was rigorously proved by Ehrenfest in 1911, namely, if one assumes statistical independence, together with Einstein’s hypothesis of quanta of light, then one should find back Wien’s law, hence a statistical behaviour of Maxwell–Boltzmann-type, rather than Planck’s law, which incorporates a statistical behaviour of Bose–Einstein type (Monaldi,, 2009; Ehrenfest,, 1911). In particular, Einstein, who was well aware of the scientific literature of the time, was the first to realize that a ‘certain type of collapse occurs in the ultraviolet regime’, as it is in this spectral region that photons primarily exhibit their quantum nature.
@@
In 1924, Bose worked out an alternative derivation of Planck’s radiation law (Bose,, 1924), and Einstein immediately followed Bose’s method, applying it to a gas of atoms or molecules (Einstein,, 1924; Einstein, 1925a, ; Einstein, 1925b, ). Einstein also observed, and this worried him, that given this lack of statistical independence, it does seem as if there is a ‘mysterious force that somehow makes the photons behave in a statistically dependent way’. These studies were appreciated by Schrödinger (Schrodinger, 1926a, ), but also marked the departure from Ehrenfest’s view of statistical mechanics. In addition, the scientific community was ready to welcome the birth of QM, and these seminal ideas of the pioneers were thus incorporated in an abstract formalism that uses the mathematics of Hilbert spaces.
@@
In modern QM, quantum mechanical statistics, namely, Bose–Einstein and Fermi–Dirac statistics, follow from two postulates, the ‘tensor product postulate’, which rules the combination, or composition, of quantum physical entities, and the ‘symmetrization postulate’, which rules the possible states of identical quantum entities, bosons and fermions. According to these postulates, identical quantum entities are ‘completely indistinguishable from a physical point of view’, and it would be this complete indistinguishability that is responsible for the lack of statistical independence identified by the pioneers. However, this formal treatment of identical quantum entities, though mathematically coherent and physically fruitful, raises long-standing epistemological and empirical issues, which are not sorted yet (see, e.g., (Dieks,, 2023) and references therein).
@@
Coming to our research, which in this article we intend to relate to the issues we have just evoked, it has moved from the above epistemological and mathematical foundations of QM to the development of a quantum mechanical framework for human cognition. More specifically, we have been interested in the mathematical modelling of how people combine and exchange words and concepts through language and how new entities of meaning are formed through the combination of concepts (Aerts & Sozzo,, 2016; Aerts, Sassoli, Sozzo,, 2016; Pisano & Sozzo,, 2020; Aerts et al.,, 2021). Indeed, growing empirical evidence has convincingly demonstrated in the last three decades that high-level cognitive processes involving perception, language, judgements and decisions, are not well represented in the mathematical formalisms originally conceived to deal with classical entities, e.g., Boolean algebras, Kolmogorovian probabilities, fuzzy sets, and complexity theory. On the contrary, the use of the mathematical formalism of quantum theory in Hilbert space has been successful as a modelling, predictive and explanatory framework to represent entities in cognitive-linguistics domains (see, e.g., (Aerts & Aerts,, 1995; Vanrijsbergen,, 2004; Aerts, 2009a, ; Pothos & Busemeyer,, 2009; Khrennikov,, 2010; Busemeyer & Bruza,, 2012; Aerts et al.,, 2013; Aerts, Gabora, Sozzo et al,, 2013; Haven & Khrennikov,, 2013; Kwamples et al.,, 2015; Dallachiara, Giuntini & Negri, 2015a, ; Dallachiara, Giuntini & Negri, 2015b, ; Melucci,, 2015; Aerts & Sozzo & Veloz,, 2016; Blutner, beim Graben,, 2016; Broekaert et al.,, 2017; Aerts et al., 2018a, ) and references therein). The reason is that quantum mechanical structures, better than classical structures, are able to cope with features such as ‘intrinsic uncertainty’, ‘contextuality’, ‘emergence’, ‘indeterminism’, and ‘superposition’, and there is growing evidence that these features are also present in human cognition, hence they are not peculiar of micro-physical entities, as is usually believed according to the standard view underlying QM.
@@
More recently, we have investigated aspects of applied cognition, e.g., computational linguistics, natural language processing and information retrieval, where the meaning content of words (thus, the corresponding concepts) and texts is recovered through digital instruments, as search engines, corpora of documents and other information retrieval systems. This investigation has allowed us to identify Bose–Einstein statistics in the combination of words, and the corresponding concepts, in texts produced by human language. We have proved through various empirical examples, that, if one considers a large text, which can be treated as a combination of several words, then this behaves at a statistical level as a ‘quantum mechanical gas of bosons’. To this end, we have attributed ‘energy levels’ to words, in an inverse proportion to their frequency of appearance in the text. Then, we have proved that the distribution of these energy levels across words does not obey Maxwell–Boltzmann statistics but, rather, Bose–Einstein statistics. We have introduced the term ‘cogniton’ as the fundamental quantum of cognition, in such a way that the meaning content of a text can be considered as a ‘quantum mechanical Bose–Einstein gas of cognitons’. Finally, we have identified a ‘mechanism of meaning dynamics’ among the words appearing in a text, which makes more frequent, thus lower energy, words to be more likely to appear in the text than less frequent words, and this behaviour closely resembles a ‘Bose–Einstein condensate at a temperature close to absolute zero’ (Aerts & Beltran,, 2020; Beltran,, 2021; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ).
@@
Recently, we have attributed the formation of entanglement, in both physics and cognitive-linguistic domains, to a ‘mechanism of contextual updating’ (Ijtp,, 2023; Philtransa,, 2023). If one writes a text, each time one adds a new word, the new word will depend on the meaning of all the others. Equivalently, each time a word is added to a text, the word ‘contextually updates the meaning content of the entire text’, because the new word not only fits meaning-wise the entire context of the text, but also gives, by being added, new meaning to the same. Later in the article, we will illustrate that this mechanism of contextual updating brings each word into a relationship of entanglement with the other words, and in turn also introduces the lack of independence of the separate words, that is, the fact that words tend to clump together, revealing a Bose–Einstein behaviour.
Returning now to physics, note that Einstein was deeply convinced that Planck’s law, not Wien’s law, was the correct one, which is why he supported Bose’s approach, although he lost at that point the support of Ehrenfest, who was instead heavily critical of Bose’s method. [^4] Ehrenfest expressed his criticism towards Bose’s method to a fellow physicist, Abram Joffé, in a letter dated 9 October 1924: “Precisely now Einstein is with us. We coincide fully with him that Bose’s disgusting work by no means can be understood in the sense that Planck’s radiation law agrees with light atoms moving independently” (see (Moskovchenko & Frenkel,, 1990), pp. 171–172). Einstein’s ideas made it possible for him to predict the phenomenon of Bose–Einstein condensation at very low temperatures, i.e. all elementary constituents tend to occupy the same micro-state, namely, the state of minimal energy. As we have mentioned, Einstein wondered what was the nature of the mysterious force at the origin of this behaviour and the corresponding violation of statistical independence of the micro-states of identical photons. Though there is currently no final answer to this question, we believe that the identification of a Bose–Einstein quantum statistics in human language might provide some valuable insights. More specifically, we consider it possible that the mysterious force that brings together identical photons can somehow be related to the physical counterpart of the ‘force of meaning’ that is active in a text that tells a story. Indeed, it can be shown, and we will do so in the remainder of this article, that the statistical dependence of micro-states, which manifests itself as a consequence of the presence of Bose–Einstein statistics, is of such a nature as to be compatible with a statistical dependence that occurs as a consequence of a mechanism of contextual updating carried by meaning.
@@
Our analysis indicates that there can be a powerful structural analogy between ‘quantum behaviour’ and ‘cognitive behaviour’. This was already clear in the development of ‘quantum cognition’, where the cognitive domain was modelled through quantum mathematics. But it is also possible to look at things the other way, noting that micro-world entities become more intelligible when interpreted as carrier of meaning (obviously, a different meaning than that of human conceptual entities). This last statement certainly would need to be extensively explored and justified, but that would be beyond the scope of this article. We simply mention here that the possibility that the microscopic world possesses a nature similar to that of the human conceptual world has been explored in depth by our group, in what we have termed the ‘conceptuality interpretation of quantum mechanics’ (Aerts, 2009b, ; Aerts et al.,, 2020).
@@
In Section 6, we establish a connection between temperature and randomness at the level of ‘words and their place’ in a text. More specifically, randomness is introduced by distributing words at random in the places they occupied in a text that tells a story. This means that we introduce the physical equivalent of the notion of ‘heat’ in our thermodynamics for human language and cognition. In this way, we draw the lines for a further elaboration of a thermodynamic theory of human language and cognition, whose details will be published in (Temperature,, 2024).
@@ 2 Identity and indistinguishability in quantum mechanics
We review in this section the mathematical representation of identical physical entities in Hilbert space, as presented in modern manuals of QM (see, e.g., (Huang,, 1987; Eisberg & Resnick,, 1985)). This formal treatment automatically incorporates the physical indistinguishability of identical entities in QM, but also raises long-standing epistemological problems, which are also briefly sketched (see, e.g., (Dieks,, 2023) and references therein). We will limit ourselves to consider the situation of two identical entities, for the sake of simplicity. However, the conclusions we reach in this section also hold in the general case of a large number of identical entities.
@@
Indeed, let us consider two identical physical entities described by classical mechanics which, at a given time, occupy two different regions of space. Then, the dynamical equations of motion guarantee that each of the two entities follows a well-defined ‘trajectory’ which allows us to distinguish them from each other. In other words, though identical, two classical physical entities can always be distinguished. The situation is radically different in QM and one can resort to various theoretical arguments, e.g., Heisenberg’s uncertainty principle which forbids the introduction of the notion of trajectory, the possible overlapping of initially separated wave functions due to the interaction between the entities, etc., and one typically maintains that there is no physical process which allows to distinguish two identical physical entities: they are ‘physically completely indistinguishable’ in QM (Huang,, 1987; Eisberg & Resnick,, 1985).
Coming to the mathematical formulation in Hilbert space, to formally deal with identical physical entities and their indistinguishability, one introduces in QM a new postulate, the ‘symmetrization postulate’, in addition to the ‘tensor product postulate’, which rules the composition of quantum entities. The symmetrization postulate requires that the state vector $|\psi(1,2)\rangle$ of two identical entities, labelled $1$ and $2$, is either ‘symmetric’ or ‘anti-symmetric’ with respect to the permutation $1\leftrightarrow 2$, that is, one has either $|\psi(2,1)\rangle=+|\psi(1,2)\rangle$ (symmetric case) or $|\psi(2,1)\rangle=-|\psi(1,2)\rangle$ (anti-symmetric case). The ‘spin-statistics theorem’, derivable within quantum field theory (Pauli,, 1940), then guarantees that physical entities with integer spin, 00, $1$, $2$, …, also called ‘bosons’, e.g., photons, are described by a symmetric state vector, while physical entities with semi-integer spin, $1/2$, $3/2$, $5/2$, …, also called ‘fermions’, e.g., electrons, are described by an anti-symmetric state vector.
@@
with $C$ and $D$ physical constants, governing the statistical behaviour of a large number of distinguishable physical entities in classical statistical mechanics (Huang,, 1987; Eisberg & Resnick,, 1985).
Quantum indistinguishability and the corresponding quantum mechanical statistics play a fundamental role in several macroscopic quantum effects which have been widely confirmed experimentally, such as chemical bonds, the properties of semi-conductors, Bose–Einstein condensation, superfluidity and superconductivity (Eisberg & Resnick,, 1985; Cornell & Wieman,, 2002; Ketterle,, 2002; Annett,, 2005).
However, the symmetrization postulate has also relevant epistemological implications. Indeed, let us limit ourselves to consider position coordinates and neglect for a moment spin variables. Let $|\psi_{A}(1)\rangle$ and $|\psi_{B}(2)\rangle$ be the unit vectors representing the states of the physical entities $1$ and $2$, and corresponding to two wave functions which are non-zero only in the widely separated regions $A$ and $B$, respectively. Then, combining Equations (1) and (2), the symmetrization postulate requires that the state vector of the composite entity made up of entities $1$ and $2$ has to be $\frac{1}{\sqrt{2}}(|\psi_{A}(1)\rangle|\psi_{B}(2)\rangle\pm|\psi_{B}(1)\rangle|\psi_{A}(2)\rangle)$. In both cases, we obtain a bi-orthogonal decomposition of a unit vector representing an entangled state, and it is well known from the standard interpretation of the coefficients of a bi-orthogonal decomposition, that in neither case one can say that one entity is localized in the space region $A$ and the other entity is localized in the space region $B$, as it would be natural in the case of well separated wave functions. According to some authors (Dieks,, 2023; Dieks & Lubberdink,, 2020), these epistemological difficulties make problematical the possibility of experimentally preparing a composite physical entity, made up of two identical quantum entities, in a pure state in which the component entities are spatially separated but entangled in spin, such as ‘EPR-Bell-type states’ (see, e.g., (Aerts & Beltran, 2022a, ; Ijtp,, 2023; Philtransa,, 2023)). Various approaches have been put forward to solve these difficulties, but there is no general consensus in the scientific community about their resolution (Dieks,, 2023; Dieks & Lubberdink,, 2020; French & Redhead,, 1988; Saunders,, 2003; Muller & Seevinck,, 2009; Krause,, 2010).

### v2

# The Origin of Quantum Mechanical Statistics: Some Insights from Research on Human Language

@@ 1 Introduction
The foundations of thermodynamics, as prescribed by classical statistical mechanics, rely on the hypothesis that, at a microscopic level, physical systems, or ‘entities’, behave (i) randomly and (ii) independently from each other [1]. A thorough analysis of the historical development of quantum mechanics (QM) reveals that the hypothesis of statistical independence started to be questioned during early quantum theory (1900–1925) and led to the birth of quantum mechanical statistics, in particular, Bose–Einstein statistics [2, 3, 4, 5, 6, 7]. However, this was done in a way that, in our opinion, is neither straightforward nor completely understood yet.
@@
Historians and philosophers of science have carefully reconstructed early quantum theory through articles and personal correspondence and generally agree that Einstein’s 1905 article on the photoelectric effect [8] can be considered as the first genuine article on the quantum nature of physical entities, more than Planck’s 1900 article on the black body radiation [9]. Indeed, Einstein was much more explicit about ‘quantization’ than Planck’s original intention, and he directly assumed the existence of what he called ‘quanta of light’, which we now call ‘photons’ [2]. At the time, Ehrenfest, a student of Boltzmann and close friend of Einstein, was in favour of his idea of quanta of light as entities in their own right, an idea that was not shared by most of the scientific community.
@@
Both Wien and Planck had followed the path of Boltzmann’s thermodynamics, at a time when the so-called ‘ultraviolet catastrophe’ of the classical theory of radiation was still not known. [^3] The expression ‘ultraviolet catastrophe’ was introduced by Ehrenfest only in 1911 [10]. However, both Einstein and Ehrenfest identified a fundamental problem in Planck’s law, namely, the latter seemed to violate the request of statistical independence of the micro-states of different photons. This, though, is a feature of the Boltzmann view of statistical mechanics, shared by all protagonists, Einstein, Ehrenfest, but also Planck, which also follows as a consequence of the Maxwell–Boltzmann distribution defining the governing statistics. The latter result was rigorously proved by Ehrenfest in 1911, namely, if one assumes statistical independence, together with Einstein’s hypothesis of quanta of light, then one should find back Wien’s law, hence a statistical behaviour of Maxwell–Boltzmann-type, rather than Planck’s law, which incorporates a statistical behaviour of Bose–Einstein type [5, 10]. In particular, Einstein, who was well aware of the scientific literature of the time, was the first to realize that a ‘certain type of collapse occurs in the ultraviolet regime’, as it is in this spectral region that photons primarily exhibit their quantum nature.
@@
In 1924, Bose worked out an alternative derivation of Planck’s radiation law [11], and Einstein immediately followed Bose’s method, applying it to a gas of atoms or molecules [12, 13, 14]. Einstein also observed, and this worried him, that given this lack of statistical independence, it does seem as if there is a ‘mysterious force that somehow makes the photons behave in a statistically dependent way’. These studies were appreciated by Schrödinger [15], but also marked the departure from Ehrenfest’s view of statistical mechanics. In addition, the scientific community was ready to welcome the birth of QM, and these seminal ideas of the pioneers were thus incorporated in an abstract formalism that uses the mathematics of Hilbert spaces.
@@
In modern QM, quantum mechanical statistics, namely, Bose–Einstein and Fermi–Dirac statistics, follow from two postulates, the ‘tensor product postulate’, which rules the combination, or composition, of quantum physical entities, and the ‘symmetrization postulate’, which rules the possible states of identical quantum entities, bosons and fermions. According to these postulates, identical quantum entities are ‘completely indistinguishable from a physical point of view’, and it would be this complete indistinguishability that is responsible for the lack of statistical independence identified by the pioneers. However, this formal treatment of identical quantum entities, though mathematically coherent and physically fruitful, raises long-standing epistemological and empirical issues, which are not sorted yet (see, e.g., [16] and references therein).
@@
Coming to our research, which in this article we intend to relate to the issues we have just evoked, it has moved from the above epistemological and mathematical foundations of QM to the development of a quantum mechanical framework for human cognition. More specifically, we have been interested in the mathematical modelling of how people combine and exchange words and concepts through language and how new entities of meaning are formed through the combination of concepts [17, 18, 19, 20]. Indeed, growing empirical evidence has convincingly demonstrated in the last three decades that high-level cognitive processes involving perception, language, judgements and decisions, are not well represented in the mathematical formalisms originally conceived to deal with classical entities, e.g., Boolean algebras, Kolmogorovian probabilities, fuzzy sets, and complexity theory. On the contrary, the use of the mathematical formalism of quantum theory in Hilbert space has been successful as a modelling, predictive and explanatory framework to represent entities in cognitive-linguistics domains (see, e.g., [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] and references therein). The reason is that quantum mechanical structures, better than classical structures, are able to cope with features such as ‘intrinsic uncertainty’, ‘contextuality’, ‘emergence’, ‘indeterminism’, and ‘superposition’, and there is growing evidence that these features are also present in human cognition, hence they are not peculiar of micro-physical entities, as is usually believed according to the standard view underlying QM.
@@
More recently, we have investigated aspects of applied cognition, e.g., computational linguistics, natural language processing and information retrieval, where the meaning content of words (thus, the corresponding concepts) and texts is recovered through digital instruments, as search engines, corpora of documents and other information retrieval systems. This investigation has allowed us to identify Bose–Einstein statistics in the combination of words, and the corresponding concepts, in texts produced by human language. We have proved through various empirical examples, that, if one considers a large text, which can be treated as a combination of several words, then this behaves at a statistical level as a ‘quantum mechanical gas of bosons’. To this end, we have attributed ‘energy levels’ to words, in an inverse proportion to their frequency of appearance in the text. Then, we have proved that the distribution of these energy levels across words does not obey Maxwell–Boltzmann statistics but, rather, Bose–Einstein statistics. We have introduced the term ‘cogniton’ as the fundamental quantum of cognition, in such a way that the meaning content of a text can be considered as a ‘quantum mechanical Bose–Einstein gas of cognitons’. Finally, we have identified a ‘mechanism of meaning dynamics’ among the words appearing in a text, which makes more frequent, thus lower energy, words to be more likely to appear in the text than less frequent words, and this behaviour closely resembles a ‘Bose–Einstein condensate at a temperature close to absolute zero’ [38, 39, 40, 41].
@@
Recently, we have attributed the formation of entanglement, in both physics and cognitive-linguistic domains, to a ‘mechanism of contextual updating’ [42, 43]. If one writes a text, each time one adds a new word, the new word will depend on the meaning of all the others. Equivalently, each time a word is added to a text, the word ‘contextually updates the meaning content of the entire text’, because the new word not only fits meaning-wise the entire context of the text, but also gives, by being added, new meaning to the same. Later in the article, we will illustrate that this mechanism of contextual updating brings each word into a relationship of entanglement with the other words, and in turn also introduces the lack of independence of the separate words, that is, the fact that words tend to clump together, revealing a Bose–Einstein behaviour.
Returning now to physics, note that Einstein was deeply convinced that Planck’s law, not Wien’s law, was the correct one, which is why he supported Bose’s approach, although he lost at that point the support of Ehrenfest, who was instead heavily critical of Bose’s method. [^4] Ehrenfest expressed his criticism towards Bose’s method to a fellow physicist, Abram Joffé, in a letter dated 9 October 1924: “Precisely now Einstein is with us. We coincide fully with him that Bose’s disgusting work by no means can be understood in the sense that Planck’s radiation law agrees with light atoms moving independently” (see [46], pp. 171–172). Einstein’s ideas made it possible for him to predict the phenomenon of Bose–Einstein condensation at very low temperatures, i.e. all elementary constituents tend to occupy the same micro-state, namely, the state of minimal energy. As we have mentioned, Einstein wondered what was the nature of the mysterious force at the origin of this behaviour and the corresponding violation of statistical independence of the micro-states of identical photons. Though there is currently no final answer to this question, we believe that the identification of a Bose–Einstein quantum statistics in human language might provide some valuable insights. More specifically, we consider it possible that the mysterious force that brings together identical photons can somehow be related to the physical counterpart of the ‘force of meaning’ that is active in a text that tells a story. Indeed, it can be shown, and we will do so in the remainder of this article, that the statistical dependence of micro-states, which manifests itself as a consequence of the presence of Bose–Einstein statistics, is of such a nature as to be compatible with a statistical dependence that occurs as a consequence of a mechanism of contextual updating carried by meaning.
@@
Our analysis indicates that there can be a powerful structural analogy between ‘quantum behaviour’ and ‘cognitive behaviour’. This was already clear in the development of ‘quantum cognition’, where the cognitive domain was modelled through quantum mathematics. But it is also possible to look at things the other way, noting that micro-world entities become more intelligible when interpreted as carrier of meaning (obviously, a different meaning than that of human conceptual entities). This last statement certainly would need to be extensively explored and justified, but that would be beyond the scope of this article. We simply mention here that the possibility that the microscopic world possesses a nature similar to that of the human conceptual world has been explored in depth by our group, in what we have termed the ‘conceptuality interpretation of quantum mechanics’ [44, 45].
@@
In Section 6, we establish a connection between temperature and randomness at the level of ‘words and their place’ in a text. More specifically, randomness is introduced by distributing words at random in the places they occupied in a text that tells a story. This means that we introduce the physical equivalent of the notion of ‘heat’ in our thermodynamics for human language and cognition. In this way, we draw the lines for a further elaboration of a thermodynamic theory of human language and cognition, whose details will be published in [47].
@@ 2 Identity and indistinguishability in quantum mechanics
We review in this section the mathematical representation of identical physical entities in Hilbert space, as presented in modern manuals of QM (see, e.g., [1, 48]). This formal treatment automatically incorporates the physical indistinguishability of identical entities in QM, but also raises long-standing epistemological problems, which are also briefly sketched (see, e.g., [16] and references therein). We will limit ourselves to consider the situation of two identical entities, for the sake of simplicity. However, the conclusions we reach in this section also hold in the general case of a large number of identical entities.
@@
Indeed, let us consider two identical physical entities described by classical mechanics which, at a given time, occupy two different regions of space. Then, the dynamical equations of motion guarantee that each of the two entities follows a well-defined ‘trajectory’ which allows us to distinguish them from each other. In other words, though identical, two classical physical entities can always be distinguished. The situation is radically different in QM and one can resort to various theoretical arguments, e.g., Heisenberg’s uncertainty principle which forbids the introduction of the notion of trajectory, the possible overlapping of initially separated wave functions due to the interaction between the entities, etc., and one typically maintains that there is no physical process which allows to distinguish two identical physical entities: they are ‘physically completely indistinguishable’ in QM [1, 48].
Coming to the mathematical formulation in Hilbert space, to formally deal with identical physical entities and their indistinguishability, one introduces in QM a new postulate, the ‘symmetrization postulate’ (also known as the ‘exchange symmetry principle’), in addition to the ‘tensor product postulate’, which rules the composition of quantum entities. The symmetrization postulate requires that the state vector $|\psi(1,2)\rangle$ of two identical entities, labelled $1$ and $2$, is either ‘symmetric’ or ‘anti-symmetric’ with respect to the permutation $1\leftrightarrow 2$, that is, one has either $|\psi(2,1)\rangle=+|\psi(1,2)\rangle$ (symmetric case) or $|\psi(2,1)\rangle=-|\psi(1,2)\rangle$ (anti-symmetric case). The ‘spin-statistics theorem’, derivable within quantum field theory [49], then guarantees that physical entities with integer spin, 00, $1$, $2$, …, also called ‘bosons’, e.g., photons, are described by a symmetric state vector, while physical entities with semi-integer spin, $1/2$, $3/2$, $5/2$, …, also called ‘fermions’, e.g., electrons, are described by an anti-symmetric state vector.
@@
with $C$ and $D$ physical constants, governing the statistical behaviour of a large number of distinguishable physical entities in classical statistical mechanics [1, 48].
Quantum indistinguishability and the corresponding quantum mechanical statistics play a fundamental role in several macroscopic quantum effects which have been widely confirmed experimentally, such as chemical bonds, the properties of semi-conductors, Bose–Einstein condensation, superfluidity and superconductivity [48, 50, 51, 52].
However, the symmetrization postulate has also relevant epistemological implications. Indeed, let us limit ourselves to consider position coordinates and neglect for a moment spin variables. Let $|\psi_{A}(1)\rangle$ and $|\psi_{B}(2)\rangle$ be the unit vectors representing the states of the physical entities $1$ and $2$, and corresponding to two wave functions which are non-zero only in the widely separated regions $A$ and $B$, respectively. Then, combining Equations (1) and (2), the symmetrization postulate requires that the state vector of the composite entity made up of entities $1$ and $2$ has to be $\frac{1}{\sqrt{2}}(|\psi_{A}(1)\rangle|\psi_{B}(2)\rangle\pm|\psi_{B}(1)\rangle|\psi_{A}(2)\rangle)$. In both cases, we obtain a bi-orthogonal decomposition of a unit vector representing an entangled state, and it is well known from the standard interpretation of the coefficients of a bi-orthogonal decomposition, that in neither case one can say that one entity is localized in the space region $A$ and the other entity is localized in the space region $B$, as it would be natural in the case of well separated wave functions. According to some authors [16, 53], these epistemological difficulties make problematical the possibility of experimentally preparing a composite physical entity, made up of two identical quantum entities, in a pure state in which the component entities are spatially separated but entangled in spin, such as ‘EPR-Bell-type states’ (see, e.g., [40, 42, 43]). Various approaches have been put forward to solve these difficulties, but there is no general consensus in the scientific community about their resolution [16, 53, 54, 55, 56, 57].